Two dice are thrown. What is the probability that the first die rolls a 1 if it is known that the number of points on the second die is higher than on the first?


1
Expert's answer
2022-04-07T08:08:46-0400

Let's A - the event of rolling a 1 of the first die.

Obviously the number of points on the second die can take on one of the values 2, 3, 4, 5, 6.

For convenience let's denote the respective events B2, B3, B4, B5, B6.

These events are equally possible, "P(B_2)=P(B_3)=P(B_4)=P(B_5)=P(B_6)=\cfrac{1}{5}."

According to the law of Total Probability,

"P(A)=P(B_2)\cdot P(A|B_2)+P(B_3)\cdot P(A|B_3)+\\
+P(B_4)\cdot P(A|B_4)+P(B_5)\cdot P(A|B_5)+\\
+P(B_6)\cdot P(A|B_6)."

Here "P(A|B_i)" are the conditional probabilities.

If there are 2 points on the second die, "P(A|B_2)=1," because the only event when the number of points on the first die is less then 2 is A.

For the event B3 there are possible values for the number of points on the first die are 1 and 2,

"P(A|B_3)=\cfrac{1}{2}."

For the event B4 there are possible values for the number of points on the first die are 1, 2 and 3,

"P(A|B_4)=\cfrac{1}{3}," and so on.

The sought probability of 1 point on the first die:

"P(A)=\cfrac{1}{5}\cdot 1+\cfrac{1}{5}\cdot \cfrac{1}{2}+\cfrac{1}{5}\cdot \cfrac{1}{3}+\cfrac{1}{5}\cdot \cfrac{1}{4}+\cfrac{1}{5}\cdot \cfrac{1}{5}=\\
=\cfrac{1}{5}\cdot \cfrac{137}{60}= \cfrac{137}{300}\approx0.457."